The Sandoval family is looking to rent a large truck for their upcoming move. With Kirk's Moving, they would pay $25 for the first day plus $2 per additional day. With Westford Rent-a-Truck, in comparison, the family would pay $23 for the first day plus $4 per additional day. Before deciding on which company to use, Mrs. Sandoval wants to find out what number of additional days would make the two choices equivalent with regards to cost. What would the total cost be? How many additional days would that be?

Answer by ikleyn(52803)

.
The Sandoval family is looking to rent a large truck for their upcoming move.
With Kirk's Moving, they would pay $25 for the first day plus $2 per additional day.
With Westford Rent-a-Truck, in comparison, the family would pay $23 for the first day plus $4 per additional day.
Before deciding on which company to use, Mrs. Sandoval wants to find out what number of additional days would make
the two choices equivalent with regards to cost.
What would the total cost be? How many additional days would that be?
~~~~~~~~~~~~~~~~~~~

With Kirk's Moving,         the cost is  25 + 2a  dollars, where "a" is "additional days".


With Westford Rent-a-Truck, the cost is  23 + 4a  dollars, where "a" is "additional days".


To find "additional days" when two companies become equivalent, write this equation

    25 + 2a = 23 + 4a


saying that two costs are the same.  From the equation


    25 - 23 = 4a - 2a 

       2    =    2a

       a    = 2/2 = 1,


ANSWER.  One additional day.  The total EQUIVALENT cost is then  25 + 2*1 = 27 dollars = 23 + 4*1.

Solved and thoroughly explained.